# Kelvin waves in nonequilibrium universal dynamics of relativistic scalar field theories

### v1

@@ IV.1 Self-similar evolution of the momentum spectrum
The O$(1)$ theory (1) in $d=2$ spatial dimensions provides a particularly intriguing and valuable framework for the study of relevant degrees of freedom for our initial conditions. Comparable simulations in $d=3$ for the massive O$(1)$ theory [31, 61] have revealed the momentum distribution function (6) which evolves self-similarly according to
@@
According to Eq. (22), the characteristic scale decreases in time as $p_{\mathrm{IR}}(t)\sim t^{-\beta}$, while the overall normalisation scales as $\mathcal{N}(t)\sim t^{2(\alpha/\kappa-\beta)}$. If particle number and thus the integral $\int\text{d}^{d}pf(t,p)$ is conserved in time, in $d$ dimensions, the relation $\alpha=d\beta$ must hold. Semiclassical simulations gave the universal scaling exponents $\beta\approx 0.5$ and $\alpha\approx 1.5\approx d\,\beta$, and $\kappa\approx 4.5$ in $d=3$ dimensions [28, 61].
@@
Figure 5: Vortex (line) density shown as a function of time, in $d=2$ and $d=3$ dimensions for O$(1)$ scalar fields. The decay reveals a power law $N_{\mathrm{v}}\sim t^{-\xi}$ with $\xi\approx 0.45$ in two dimensions, and $\xi\approx 0.91$ in three dimensions.
@@ IV.2 Coarsening dynamics
Based on the spectral and statistical functions, the above excitation dynamics for single-component scalar fields was found to be markedly different from what is observed and expected in the large-$N$ limit [31], where we have an analytic understanding of these exponents [28, 29, 30]. On the other hand, the exponent $\beta\approx 0.5$ is known to apply in qualitatively very different situations. As an alternative to the self-similar transport of free Goldstone quasiparticle excitations to lower momenta, which dominates the large-$N$ limit [30, 81], it could also result from diffusion-type coarsening dynamics of the spatial field configuration, e.g., as induced by mutual annihilation of vortices and antivortices excited in the system [74].
Yet, in two-dimensional O$(1)$ simulations, the equal-time distribution function has only been found to scale with $\beta\approx 0.25$, $\alpha\approx 0.5$, and $\kappa\approx 4$ [76]. This is confirmed by our results in Fig. 4a, where we show the evolution of $f(t,p)$ for a two-dimensional system in the main panel and its self-similar scaling in the inset using the stated exponents. Despite scanning through a wide range of initial conditions, these results remained unchanged. In particular, we were unable to find a stable $\beta\approx 0.5$ scaling regime in $d=2$. In contrast, for $N>1$, we have easily recovered the diffusion-type scaling with $\beta\approx 0.5$ and $\alpha=d\beta\approx 1.0$.
@@
A closely related scaling exponent to the $d=2$ case of $\mathrm{O}(1)$ theory, $\beta\approx 0.2$ has been found in simulations of a nonrelativistic Bose gas in two dimensions upon imprinting vortices into the initial field configuration [74], also confirmed experimentally [11]. This scaling behaviour is phenomenologically understood as due to three-body collisions of bound vortex-antivortex pairs with nearby vortices or antivortices. Any such collision allows for an exchange of vortices after which the newly formed vortex-antivortex pair can be more strongly bound and thus potentially undergo fast subsequent vortex-antivortex annihilation [74]. Hence, pair annihilation is dominated by the three-body collision rate, which gives rise to the strongly subdiffusive exponent $\beta\approx 0.2\ll 0.5$. This stands in contrast to isolated vortex-antivortex annihilation due to background fluctuations, which is governed by a diffusion-type scaling with $\beta\approx 0.5$ [74, 78, 11, 91]. The relevance of three-body collisions was likewise identified in Ref. [76] for the $\mathrm{O}(1)$ model in two dimensions without imprinting vortices.
In both Refs. [74, 76] the average vortex density was found to decay as a power law, $n_{\mathrm{v}}\sim t^{-\xi}$ with $\xi\approx 0.41\approx 2\beta$, which reflects that the decreasing infrared scale $p_{\mathrm{IR}}(t)\sim t^{-\beta}$ corresponds to a growing characteristic length scale, which can be associated with the mean inter-defect distance $\ell_{\mathrm{v}}\sim p_{\mathrm{IR}}^{-1}\sim t^{\beta}$. Hence, it is useful to define the exponent
@@
An interesting question therefore is to what extent the spatiotemporal scaling of $f(t,p)$ and the momentum power law fall-off is, in general, driven by coarsening, or whether the two scaling phenomena are independent. After all, the value of the exponent $\kappa$ and thus the steepness of the distribution determines whether and how particles can be transported towards the infrared and eventually allow condensation in the $p=0$ mode. The buildup of such a transport process necessarily involves some underlying coarsening due to causality, as explained in Sec. II. As indicated by the numerical simulations in [81], different processes driven by either, e.g., topological defects, or compressible sound excitations, can dominate the overall scaling evolution.
Hence, before moving on to investigating the excitation spectrum from unequal-time correlation functions, we examine the dynamics of vortices in two dimensions and vortex lines in three dimensions explicitly and extract the respective $\beta_{\mathrm{L}}$ exponents from the spatial field configurations. As mentioned in Sec. III, (anti)vortex defects are identified by a ±$\pm 2\pi$ phase winding in $\theta(t,\mathbf{x})$, which coincide with a depression in the energy density on the lattice. Based on this, we have extracted the total vortex (line) density in two and three dimensions, which we show as functions of time in Fig. 5. For additional details on the data analysis and the extraction of these quantities, see App. A.
This vortex density reveals a power law $\sim t^{-\xi}$ with $\xi\approx 0.45$ or $\beta_{\mathrm{L}}\approx 0.23$ for the $\mathrm{O}(1)$ model in two dimensions, which is consistent with the $\beta\approx 0.25$ self-similar scaling of the equal-time distribution function $f$.  [^3] Note however that, in two dimensions, vortex coarsening could also be characterised by $\beta_{\mathrm{L}}\approx\beta\approx 0.5$, as seen for a nonrelativistic Bose gas in [74] and discussed above. In three dimensions, we find $\xi\approx 0.91$. It represents the decrease of the total vortex line length per unit volume, which corresponds to a dilution of the vortex rings and tangles as $\sim\ell_{\text{v}}^{-2}\sim t^{-2\beta}$ [156, 71], where $\ell_{\text{v}}$ is the scale measuring the mean distance between vortex lines. Hence, also in $d=3$, one has $\xi=2\beta$, and the decrease of the vortex line density implies $\beta_{\mathrm{L}}\approx 0.46\approx\beta$. While these numbers look convincing, a clearer link to the relevant quasiparticle excitations in $f$ would be desirable. To make such a connection, in the next subsection we discuss the extraction of the frequency and momentum dependence of the correlation functions that are closely related to the distribution function (6) but additionally encode the systems’ excitation spectra.
@@ IV.3 Unequal-time correlation functions
For both the two- and three-dimensional theories, the statistical function $F$ is dominated by a single peak at low momenta in the infrared, which is accurately captured by the (non-Lorentzian) hyperbolic secant function, [^4] Regarding the notation: this peak corresponds to the “transport peak” in [122] and the “non-Lorentzian peak” (nL) in [31].
@@
Example data extracted from numerical simulations for $F(t,\omega,p)$ and $\rho(t,\omega,p)$ is shown in Fig. 6 for two dimensions, as functions of frequency $\omega-M$ relative to the ‘mass’ gap of the dispersion, $\omega(p=0)=M$, for different times $t$ and different momenta $p$. To obtain a clearer picture, it is useful to depict the statistical function $F(t,\omega,p)$ in the frequency-momentum plane, as shown in Fig. 7, in both two (left) and three (right) dimensions, evaluated at time $t=1000$. As already denoted on these figures and explained in the following section, the $F_{\mathrm{K}}$ peak is identified with Kelvin waves and their quantised excitations, kelvons, in three dimensions, as well as the analogous kelvon quasiparticles in two dimensions. From both representations, we can also see that at higher momenta, a second peak appears, which has been identified before as a Bogoliubov peak [31]. This Bogoliubov peak is more easily accessible from the spectral function $\rho(t,\omega,p)$, which is what we have used at $p=0$ to determine the effective mass gap $M$. The peak can be fitted by approximate Lorentzians of the form
@@
that capture the positive and negative branches of the Bogoliubov excitation. Moreover, its dispersion can be accurately fitted by using the relativistic generalisation (18) of the Bogoliubov dispersion [31], which is shown in Fig. 8.
@@
We note that in general, each of the excitations in $F$ should have a counterpart in $\rho$. For the Bogoliubov peak, we confirm this for the two-dimensional theory in Fig. 6 and refer to Ref. [31] for a confirmation in $d=3$. In contrast, the identification of the $F_{\mathrm{K}}$ peak given by (25) in $\rho$ is challenging due to the peak’s relative suppression in comparison to the Bogoliubov excitations. For $d=3$, the existence of such a peak in $\rho$ was demonstrated in Ref. [31] by carefully analysing high-statistics data. For $d=2$, our level of statistics is insufficient to give a clear statement about its existence in $\rho$.
@@ IV.4 Kelvon dispersion relations
In three dimensions, the dispersion relation of this peak has been extracted before and was considered to be approximately linear [122, 31]. However, upon closer inspection, one can see that it is, in fact, quadratic at low momenta and linear at higher momenta, which is characteristic for kelvons [154]. Re-analysing the dispersion relation from Ref. [31], we find an excellent fit with the Kelvin wave dispersions as shown in Fig. 9b. We fitted the long-wavelength limit of Eq. (20), with the additional zero-momentum shift proposed in Ref. [152],
@@
where $\omega_{0}$ is the frequency of the kelvon with the smallest momentum $p_{0}\to 0$ relative to the mass gap $M$ extracted from Bogoliubov excitations, $\Gamma$ is the (here positive) vortex circulation and $r_{\mathrm{c}}$ is the vortex core radius. These quantities show a time dependence, which represents characteristics of the scaling dynamics not extracted previously. We find an increasing vortex core radius $r_{c}$, which reflects that in the course of the infrared transport, the number of coherent excitations at momenta on the order of the inverse healing length, [^5] The healing length was extracted from the linear part of the Bogoliubov dispersion, where the speed of sound $c_{\mathrm{s}}$ and the healing length $p_{\xi}$ are related via $c_{\mathrm{s}}=p_{\xi}/\sqrt{2}M$. $p_{\xi}=\xi_{\text{h}}^{-1}\approx 1.2\,Q$ in $d=2$ and $p_{\xi}\approx 0.6\,Q$ in $d=3$, decreases in time. This increase in the core radius as $r_{c}(t)\sim t^{\beta}$ is corroborated by the findings of [71], where the total volume enclosed by the vortex cores in decaying Vinen turbulence in a superfluid in $d=3$ dimensions was found to settle to a constant at large times, while the vortex line density decays as $t^{-1}$. The gap $\omega_{0}$, in $d=3$, is found to be negative and its modulus to decrease as $|\omega_{0}(t)|\sim t^{-2\beta}$, as can be observed in Fig. 9(b). This decrease is consistent with the expectation that $\omega_{0}$ corresponds to the energy of the Kelvin mode with the lowest wave number along the vortex lines, which should decrease as $p_{0}^{2}(t)\sim t^{-2\beta}$ for vortex rings with diameter increasing as $\ell_{\text{v}}(t)\sim t^{\beta}$. As a result, the dispersions are found to obey the self-similar scaling relation
@@
with exponent $\beta\approx 0.5$ and $z=2$ [31], as also visible in the inset. Here, $z=2$ reflects the dynamical scaling dimension of the Kelvin-wave dispersion relation. At later times, we fit the dispersion (27) across a progressively smaller momentum range, as shown by the dashed lines in Fig. 9(b). This is necessary because the dispersion gradually starts to flatten at higher momenta, as clearly visible at $t=4000$, deviating from the expected form and making the fit increasingly unreliable. As a result, the extracted parameters become less meaningful in this regime. A more detailed discussion of this behaviour can be found in App. B.
For two dimensions, we show the dispersion relation at different times in Fig. 9a. We find that it can be equally well fitted by both the kelvon and linear Bogoliubov dispersions, which we discuss more quantitatively in App. C by investigating the residuals for both $d=2,3$ and both fits. The dispersion obeys the same self-similar scaling form (28) as in $d=3$, but with exponents $\beta\approx 0.25$ and again $z\approx 2$, as shown in the inset of Fig. 9a.
@@
One of the prominent differences between two and three dimensions is that the peaks are considerably broader in two dimensions, as observed from the frequency-momentum resolved statistical function $F(t,\omega,p)$ in Fig. 7. A similar effect occurs in non-Abelian gauge plasmas where the excitation spectrum in two dimensions exhibits a broad non-Lorentzian peak that remains broad even at late times [157] while the quasiparticle peak in three-dimensional plasmas becomes narrower with time [141]. In general, fluctuations are more significant in lower dimensions because of the reduced spatial volume available to average them out. This leads to more prominent noise and dissipation effects, which can then further broaden the peaks.
Additionally, the broadening in two dimensions might also reflect the stronger coupling between the different modes – kelvons and phonons from vortex interactions – due to the reduced dimensionality, where interactions and energy exchange between excitations are more pronounced. This interplay could obscure the distinction between different quasiparticle contributions, making it more challenging to isolate specific dispersion features as compared with the three-dimensional case. While in three dimensions, a distinct quadratic bending at the lowest momenta can be observed in Fig. 7b, extracting $F(t,\omega_{\mathrm{max}},p)$ as shown in Fig. 9b clarifies that, indeed, the dispersion is accurately described by a Kelvin-wave dispersion. Moreover, this three-dimensional representation also shows a striking similarity to the “space-time resolved mass spectrum” in Ref. [154], which used GPE simulations to study Kelvin and (Bogoliubov) sound waves.
Curiously, kelvons seem to dominate the dynamics of the infrared peak in $F(t,\omega,p)$, and we do not find any significant contribution from phonons arising, e.g., from vortex-ring annihilation. In two dimensions, this is likely the case as well, but the broader peaks make it more difficult to distinguish their contributions with certainty, leaving open the possibility that phonons may still play a role. How phonon generation occurs in vortex decay in two-dimensional superfluids is still a matter of active research, with different $n$-body processes put forward for the mechanisms [74, 158, 159, 78, 160]. We note that in our case, the process in two dimensions seems to be dominated by three-body collisions, where a vortex-antivortex pair and a lone vortex/antivortex are needed, as described in Sec. IV.2 and Refs. [74, 76].

### v2

@@ IV.1 Self-similar evolution of the momentum spectrum
The O(1)(1) theory (1) in $d=2$ spatial dimensions provides a particularly intriguing and valuable framework for the study of relevant degrees of freedom for our initial conditions. Comparable simulations in $d=3$ for the massive O(1)(1) theory Boguslavski and Piñeiro Orioli (2020); Moore (2016) have revealed the momentum distribution function (6) which evolves self-similarly according to
@@
According to Eq. (22), the characteristic scale decreases in time as $p_{\mathrm{IR}}(t)\sim t^{-\beta}$, while the overall normalisation scales as $\mathcal{N}(t)\sim t^{2(\alpha/\kappa-\beta)}$. If particle number and thus the integral $\int\text{d}^{d}pf(t,p)$ is conserved in time, in $d$ dimensions, the relation $\alpha=d\beta$ must hold. Semiclassical simulations gave the universal scaling exponents $\beta\approx 0.5$ and $\alpha\approx 1.5\approx d\,\beta$, and $\kappa\approx 4.5$ in $d=3$ dimensions Piñeiro Orioli et al. (2015); Moore (2016).
@@
Figure 5: Vortex (line) density shown as a function of time, in $d=2$ and $d=3$ dimensions for O(1)(1) scalar fields. The decay reveals a power law $N_{\mathrm{v}}\sim t^{-\xi}$ with $\xi\approx 0.45$ in two dimensions, and $\xi\approx 0.91$ in three dimensions.
@@ IV.2 Coarsening dynamics
Based on the spectral and statistical functions, the above excitation dynamics for single-component scalar fields was found to be markedly different from what is observed and expected in the large-$N$ limit Boguslavski and Piñeiro Orioli (2020), where we have an analytic understanding of these exponents Piñeiro Orioli et al. (2015); Walz et al. (2018); Chantesana et al. (2019). On the other hand, the exponent $\beta\approx 0.5$ is known to apply in qualitatively very different situations. As an alternative to the self-similar transport of free Goldstone quasiparticle excitations to lower momenta, which dominates the large-$N$ limit Chantesana et al. (2019); Mikheev et al. (2019), it could also result from diffusion-type coarsening dynamics of the spatial field configuration, e.g., as induced by mutual annihilation of vortices and antivortices excited in the system Karl and Gasenzer (2017).
Yet, in two-dimensional O(1)(1) simulations, the equal-time distribution function has only been found to scale with $\beta\approx 0.25$, $\alpha\approx 0.5$, and $\kappa\approx 4$ Deng et al. (2018). The temporal scaling exponents $\alpha$ and $\beta$ are confirmed by our results in Fig. 4a, where we show the evolution of $f(t,p)$ for a two-dimensional system in the main panel and its self-similar scaling in the inset using the stated exponents. We find the power law decrease of the distribution function to be closer to $\sim p^{-3.5}$ for the relatively short momentum interval shown, and we do not exclude the previously observed value of $\kappa$. Despite scanning through a wide range of initial conditions, these results remained unchanged. In particular, we were unable to find a stable $\beta\approx 0.5$ scaling regime in $d=2$. In contrast, for $N>1$, we have easily recovered the diffusion-type scaling with $\beta\approx 0.5$ and $\alpha=d\beta\approx 1.0$.
@@
A closely related scaling exponent to the $d=2$ case of $\mathrm{O}(1)$ theory, $\beta\approx 0.2$ has been found in simulations of a nonrelativistic Bose gas in two dimensions upon imprinting vortices into the initial field configuration Karl and Gasenzer (2017), also confirmed experimentally Johnstone et al. (2019). This scaling behaviour is phenomenologically understood as due to three-body collisions of bound vortex-antivortex pairs with nearby vortices or antivortices. Any such collision allows for an exchange of vortices after which the newly formed vortex-antivortex pair can be more strongly bound and thus potentially undergo fast subsequent vortex-antivortex annihilation Karl and Gasenzer (2017). Hence, pair annihilation is dominated by the three-body collision rate, which gives rise to the strongly subdiffusive exponent $\beta\approx 0.2\ll 0.5$. This stands in contrast to isolated vortex-antivortex annihilation due to background fluctuations, which is governed by a diffusion-type scaling with $\beta\approx 0.5$ Karl and Gasenzer (2017); Baggaley and Barenghi (2018); Johnstone et al. (2019); Groszek et al. (2021). The relevance of three-body collisions was likewise identified in Ref. Deng et al. (2018) for the $\mathrm{O}(1)$ model in two dimensions without imprinting vortices.
In both Refs. Karl and Gasenzer (2017); Deng et al. (2018) the average vortex density was found to decay as a power law, $n_{\mathrm{v}}\sim t^{-\xi}$ with $\xi\approx 0.41\approx 2\beta$, which reflects that the decreasing infrared scale $p_{\mathrm{IR}}(t)\sim t^{-\beta}$ corresponds to a growing characteristic length scale, which can be associated with the mean inter-defect distance $\ell_{\mathrm{v}}\sim p_{\mathrm{IR}}^{-1}\sim t^{\beta}$. Hence, it is useful to define the exponent
@@
An interesting question therefore is to what extent the spatiotemporal scaling of $f(t,p)$ and the momentum power law fall-off is, in general, driven by coarsening, or whether the two scaling phenomena are independent. After all, the value of the exponent $\kappa$ and thus the steepness of the distribution determines whether and how particles can be transported towards the infrared and eventually allow condensation in the $p=0$ mode. The buildup of such a transport process necessarily involves some underlying coarsening due to causality, as explained in Sec. II. As indicated by the numerical simulations in Mikheev et al. (2019), different processes driven by either, e.g., topological defects, or compressible sound excitations, can dominate the overall scaling evolution.
Hence, before moving on to investigating the excitation spectrum from unequal-time correlation functions, we examine the dynamics of vortices in two dimensions and vortex lines in three dimensions explicitly and extract the respective $\beta_{\mathrm{L}}$ exponents from the spatial field configurations. As mentioned in Sec. III, (anti)vortex defects are identified by a $\pm 2\pi$ phase winding in $\theta(t,\mathbf{x})$, which coincide with a depression in the energy density on the lattice. Based on this, we have extracted the total vortex (line) density in two and three dimensions, which we show as functions of time in Fig. 5. For additional details on the data analysis and the extraction of these quantities, see App. A.
This vortex density reveals a power law $\sim t^{-\xi}$ with $\xi\approx 0.45$ or $\beta_{\mathrm{L}}\approx 0.23$ for the $\mathrm{O}(1)$ model in two dimensions, which is consistent with the $\beta\approx 0.25$ self-similar scaling of the equal-time distribution function ff. [^3] Note however that, in two dimensions, vortex coarsening could also be characterised by $\beta_{\mathrm{L}}\approx\beta\approx 0.5$, as seen for a nonrelativistic Bose gas in Karl and Gasenzer (2017) and discussed above. In three dimensions, we find $\xi\approx 0.91$. It represents the decrease of the total vortex line length per unit volume, which corresponds to a dilution of the vortex rings and tangles as $\sim\ell_{\text{v}}^{-2}\sim t^{-2\beta}$ Vinen (1957); Villois et al. (2016), where $\ell_{\text{v}}$ is the scale measuring the mean distance between vortex lines. Hence, also in $d=3$, one has $\xi=2\beta$, and the decrease of the vortex line density implies $\beta_{\mathrm{L}}\approx 0.46\approx\beta$. While these numbers look convincing, a clearer link to the relevant quasiparticle excitations in $f$ would be desirable. To make such a connection, in the next subsection we discuss the extraction of the frequency and momentum dependence of the correlation functions that are closely related to the distribution function (6) but additionally encode the systems’ excitation spectra.
@@ IV.3 Unequal-time correlation functions
For both the two- and three-dimensional theories, the statistical function $F$ is dominated by a single peak at low momenta in the infrared, which is accurately captured by the (non-Lorentzian) hyperbolic secant function, [^4] Regarding the notation: this peak corresponds to the “transport peak” in Piñeiro Orioli and Berges (2019) and the “non-Lorentzian peak” (nL) in Boguslavski and Piñeiro Orioli (2020).
@@
Example data extracted from numerical simulations for $F(t,\omega,p)$ and $\rho(t,\omega,p)$ is shown in Fig. 6 for two dimensions, as functions of frequency $\omega-M$ relative to the ‘mass’ gap of the dispersion, $\omega(p=0)=M$, for different times $t$ and different momenta $p$. To obtain a clearer picture, it is useful to depict the statistical function $F(t,\omega,p)$ in the frequency-momentum plane, as shown in Fig. 7, in both two (left) and three (right) dimensions, evaluated at time $t=1000$. As already denoted on these figures and explained in the following section, the $F_{\mathrm{K}}$ peak is identified with Kelvin waves and their quantised excitations, kelvons, in three dimensions, as well as the analogous kelvon quasiparticles in two dimensions. From both representations, we can also see that at higher momenta, a second peak appears, which has been identified before as a Bogoliubov peak Boguslavski and Piñeiro Orioli (2020). This Bogoliubov peak is more easily accessible from the spectral function $\rho(t,\omega,p)$, which is what we have used at $p=0$ to determine the effective mass gap $M$. The peak can be fitted by approximate Lorentzians of the form
@@
that capture the positive and negative branches of the Bogoliubov excitation. Moreover, its dispersion can be accurately fitted by using the relativistic generalisation (18) of the Bogoliubov dispersion Boguslavski and Piñeiro Orioli (2020), which is shown in Fig. 8.
@@
We note that in general, each of the excitations in $F$ should have a counterpart in $\rho$. For the Bogoliubov peak, we confirm this for the two-dimensional theory in Fig. 6 and refer to Ref. Boguslavski and Piñeiro Orioli (2020) for a confirmation in $d=3$. In contrast, the identification of the $F_{\mathrm{K}}$ peak given by (25) in $\rho$ is challenging due to the peak’s relative suppression in comparison to the Bogoliubov excitations. For $d=3$, the existence of such a peak in $\rho$ was demonstrated in Ref. Boguslavski and Piñeiro Orioli (2020) by carefully analysing high-statistics data. For $d=2$, our level of statistics is insufficient to give a clear statement about its existence in $\rho$.
@@ IV.4 Kelvon dispersion relations
In three dimensions, the dispersion relation of this peak has been extracted before and was considered to be approximately linear Piñeiro Orioli and Berges (2019); Boguslavski and Piñeiro Orioli (2020). However, upon closer inspection, one can see that it is, in fact, quadratic at low momenta and linear at higher momenta, which is characteristic for kelvons Clark Di Leoni et al. (2015). Re-analysing the dispersion relation from Ref. Boguslavski and Piñeiro Orioli (2020), we find an excellent fit with the Kelvin wave dispersions as shown in Fig. 9b. We fitted the long-wavelength limit of Eq. (20), with the additional zero-momentum shift proposed in Ref. Simula et al. (2008b),
@@
where $\omega_{0}$ is the frequency of the kelvon with the smallest momentum $p_{0}\to 0$ relative to the mass gap $M$ extracted from Bogoliubov excitations, $\Gamma$ is the (here positive) vortex circulation and $r_{\mathrm{c}}$ is the vortex core radius. These quantities show a time dependence, which represents characteristics of the scaling dynamics not extracted previously. We find an increasing vortex core radius $r_{c}$, which reflects that in the course of the infrared transport, the number of coherent excitations at momenta on the order of the inverse healing length, [^5] The healing length was extracted from the linear part of the Bogoliubov dispersion, where the speed of sound $c_{\mathrm{s}}$ and the healing length $p_{\xi}$ are related via $c_{\mathrm{s}}=p_{\xi}/\sqrt{2}M$. $p_{\xi}=\xi_{\text{h}}^{-1}\approx 1.2\,Q$ in $d=2$ and $p_{\xi}\approx 0.6\,Q$ in $d=3$, decreases in time. This increase in the core radius as $r_{c}(t)\sim t^{\beta}$ is corroborated by the findings of Villois et al. (2016), where the total volume enclosed by the vortex cores in decaying Vinen turbulence in a superfluid in $d=3$ dimensions was found to settle to a constant at large times, while the vortex line density decays as $t^{-1}$. The gap $\omega_{0}$, in $d=3$, is found to be negative and its modulus to decrease as $|\omega_{0}(t)|\sim t^{-2\beta}$, as can be observed in Fig. 9(b). This decrease is consistent with the expectation that $\omega_{0}$ corresponds to the energy of the Kelvin mode with the lowest wave number along the vortex lines, which should decrease as $p_{0}^{2}(t)\sim t^{-2\beta}$ for vortex rings with diameter increasing as $\ell_{\text{v}}(t)\sim t^{\beta}$. As a result, the dispersions are found to obey the self-similar scaling relation
@@
with exponent $\beta\approx 0.5$ and $z=2$ Boguslavski and Piñeiro Orioli (2020), as also visible in the inset. Here, $z=2$ reflects the dynamical scaling dimension of the Kelvin-wave dispersion relation. At later times, we fit the dispersion (27) across a progressively smaller momentum range, as shown by the dashed lines in Fig. 9(b). This is necessary because the dispersion gradually starts to flatten at higher momenta, as clearly visible at $t=4000$, deviating from the expected form and making the fit increasingly unreliable. As a result, the extracted parameters become less meaningful in this regime. A more detailed discussion of this behaviour can be found in App. B.
For two dimensions, we show the dispersion relation at different times in Fig. 9a. We find that it can be equally well fitted by both the kelvon and linear Bogoliubov dispersions, which we discuss more quantitatively in App. C by investigating the residuals for both $d=2,3$ and both fits. The dispersion obeys the self-similar scaling form (28) as in $d=3$, but with exponents $\beta\approx 0.25$ and again $z\approx 2$, cf.  inset of Fig. 9a.
@@
One of the prominent differences between two and three dimensions is that the peaks are considerably broader in two dimensions, as observed from the frequency-momentum resolved statistical function $F(t,\omega,p)$ in Fig. 7. A similar effect occurs in non-Abelian gauge plasmas where the excitation spectrum in two dimensions exhibits a broad non-Lorentzian peak that remains broad even at late times Boguslavski et al. (2021) while the quasiparticle peak in three-dimensional plasmas becomes narrower with time Boguslavski et al. (2018). In general, fluctuations are more significant in lower dimensions because of the reduced spatial volume available to average them out. This leads to more prominent noise and dissipation effects, which can then further broaden the peaks.
Additionally, the broadening in two dimensions might also reflect the stronger coupling between the different modes – kelvons and phonons from vortex interactions – due to the reduced dimensionality, where interactions and energy exchange between excitations are more pronounced. This interplay could obscure the distinction between different quasiparticle contributions, making it more challenging to isolate specific dispersion features as compared with the three-dimensional case. While in three dimensions, a distinct quadratic bending at the lowest momenta can be observed in Fig. 7b, extracting $F(t,\omega_{\mathrm{max}},p)$ as shown in Fig. 9b clarifies that, indeed, the dispersion is accurately described by a Kelvin-wave dispersion. Moreover, this three-dimensional representation also shows a striking similarity to the “space-time resolved mass spectrum” in Ref. Clark Di Leoni et al. (2015), which used GPE simulations to study Kelvin and (Bogoliubov) sound waves.
Curiously, kelvons seem to dominate the dynamics of the infrared peak in $F(t,\omega,p)$, and we do not find any significant contribution from phonons arising, e.g., from vortex-ring annihilation. In two dimensions, this is likely the case as well, but the broader peaks make it more difficult to distinguish their contributions with certainty, leaving open the possibility that phonons may still play a role. How phonon generation occurs in vortex decay in two-dimensional superfluids is still a matter of active research, with different $n$-body processes put forward for the mechanisms Karl and Gasenzer (2017); Groszek et al. (2016); Yu et al. (2016); Baggaley and Barenghi (2018); Kanai and Zhang (2024). We note that in our case, the process in two dimensions seems to be dominated by three-body collisions, where a vortex-antivortex pair and a lone vortex/antivortex are needed, as described in Sec. IV.2 and Refs. Karl and Gasenzer (2017); Deng et al. (2018).
@@
We finally remark that, despite the smooth dependence of the dispersions on momentum, as seen in Figs. 79, as well as the correspondingly smooth occupation number spectra shown in Fig. 4, which are dominated by the kelvon transport peak at low momenta, cf. Fig. 6, the different Kelvon modes generically contribute to the overall field pattern in an incoherent manner. Hence, the macroscopic occupations at low momenta give rise to a kind of quasicondensate, with similar occupancies spread over a wide range of momenta, but rather random relative phases between the modes, cf. Mikheev et al. (2019) for the case of sound modes. This is reminiscent of quasicondensates in one- and two-dimensional superfluids in equilibrium, at zero and non-zero temperatures, and gives rise to the chaotic displacement excitations of the vortices seen in Fig. 2.